OpenStudy (anonymous):

PLEASE HELP! DON'T KNOW HOW TO DO THIS!! Find the slope of the line that passes through each pair of points. (-2, -8) and (1, 4)

OpenStudy (anonymous):

is there a link

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (texaschic101):

hartnn is correct... slope(m) = (y2 - y1) / (x2 - x1) (-2,-8)..x1 = -2 and y1 = -8 (1,4)..x2 = 1 and y2 = 4 now just sub slope(m) = (4 - (-8) / (1 - (-2) slope(m) = (4 + 8) / (1 + 2) can you finish ?

OpenStudy (anonymous):

sorry did not see this until now

OpenStudy (marinedude42):

thats how you solve

OpenStudy (anonymous):

I can't really understand this can someone tell me in steps? sorry I am having trouble understanding

OpenStudy (anonymous):

I finally understand this Thank you! :)